# Short-distance thermal phase structure of charged black holes in 4D Einstein–Gauss–Bonnet gravity

### v4

@@ Abstract
Glavan and Lin’s proposal of an effective four-dimensional Einstein–Gauss–Bonnet (4D-EGB) gravity framework yields predictions that differ from general relativity in some regimes. A range of black hole studies have offered insights into the dynamical and phenomenological aspects of this effective theory of gravity. In this work, we study thermodynamics of a charged 4D-EGB black hole with Gauss–Bonnet (GB) coupling $\alpha$, characterized by mass $M$ and charge $Q$ in the non-extremal regime $M>\sqrt{Q^{2}+\alpha}$ by combining a non-perturbative, quantum-gravity-inspired exponential correction to the entropy (quantified by $\eta$) with information-geometric diagnostics. Working in a canonical ensemble (fixed $Q$) paradigm, we identify thermodynamic stability regions and phase-transition-like features as the black hole size tends toward extremality due to Hawking evaporation. We then construct the Ruppeiner metric on the $(M,Q)$ state space and evaluate the associated thermodynamic curvature to characterize the effective interaction signatures and its relation to critical behavior. In addition, an effective quantum-work quantity, defined from the free-energy landscape using Jarzynski equality, is evaluated as an additional probe of short-distance, near-extremal behavior. The results indicate that departures from the general-relativistic behavior are negligible for large black holes but can become relevant at small horizon scales. Specifically, on short-distance scales, the combined influence of $\alpha$ and $\eta$ can modify stability of the extremal black hole geometry and remnants within this thermodynamic model.
@@ I Introduction
Black hole thermodynamics originates from the work of Hawking Hawking (1974, 1975) and Bekenstein Bekenstein (1973) in semiclassical gravity, where entropy satisfies the area law,
@@
with $A$ the horizon area and $\ell_{\rm p}$ the Planck length. [^1] In this paper, the natural units $c=G=\hbar=1$ are used throughout. The area scaling is closely related to holography Susskind (1995); Bousso (2003). While this framework provides a consistent macroscopic description in terms of conserved charges, the microscopic origin of black hole entropy and the statistical interpretation of thermodynamic response functions remain open Davies (1978); Page (2005).
@@
Hawking evaporation Hawking (1974, 1975) is derived in a semiclassical regime, and motivates corrections to the Bekenstein–Hawking description when the black hole approaches microscopic scales. Many quantum-gravity approaches imply a minimum length and associated departures from classical geometry Hossenfelder (2013), which in turn suggest modifications of the entropy Mann and Solodukhin (1998); Upadhyay et al. (2018); Pourhassan et al. (2020) and connections to microstate counting Strominger and Vafa (1996). Entropy corrections predicted in various frameworks include perturbative (logarithmic or power-law) terms Rovelli (1996); Ashtekar et al. (1998); Kaul and Majumdar (2000); Dabholkar et al. (2011); Mandal and Sen (2010) and non-perturbative contributions of exponential type Ashtekar (1991); Ghosh and Mitra (2014); Dabholkar et al. (2015); Chatterjee and Ghosh (2020). Examples of the latter arise in AdS/CFT-based analyses Maldacena (1998) and in treatments involving Kloosterman sums Murthy and Pioline (2009); Dabholkar et al. (2013, 2015). Such corrections are suppressed for macroscopic horizons but can dominate the thermodynamics near microscopic scales Dabholkar et al. (2013, 2015); Chatterjee and Ghosh (2020); see Refs. Hemming and Thorlacius (2007); Gregory et al. (2008); Rocha (2008); Saraswat and Afshordi (2020); Mann and Solodukhin (1998); Sen (2012); Govindarajan et al. (2001); Pourhassan et al. (2017, 2018, 2020, 2021a); Upadhyay et al. (2022); Ghaffarnejad and Ghasemi (2022); Biswas et al. (2021); Pourhassan et al. (2022a, b); Aounallah et al. (2022) for broader discussions.
@@
A complete quantum-gravity treatment would modify both geometry and thermodynamics of black holes. Nevertheless, it is common in phenomenological studies to implement quantum effects through an effective entropy deformation while keeping the classical background fixed, thereby isolating their impact on response functions and stability criteria Masood A. S. Bukhari et al. (2023); Pourhassan et al. (2022a, b); Iso and Okazawa (2011). Since thermodynamic behavior depends on the underlying gravitational dynamics, it is also of interest to examine these effects in modified-gravity models.
@@
General relativity (GR) is consistent with a wide range of observations, including gravitational waves Abbott et al. (2016a, b) and black hole shadow measurements Akiyama et al. (2019), but it also motivates extensions in regimes where curvature becomes large. Higher-curvature theories provide a concrete class of such extensions Capozziello and De Laurentis (2011); Berti et al. (2015); Shankaranarayanan and Johnson (2022). Gauss-Bonnet (GB) gravity, originating with Lanczos and Lovelock Lanczos (1938); Lovelock (1971, 1972), yields second-order field equations and avoids Ostrogradsky-type instabilities Ostrogradsky (1850), and quadratic curvature terms appear naturally in string-inspired effective actions Zwiebach (1985); Gross and Witten (1986); Gross and Sloan (1987). In four dimensions the GB invariant does not contribute to the classical field equations unless coupled to additional fields, such as scalars Blázquez-Salcedo et al. (2016); Konoplya et al. (2020); Maselli et al. (2015); Ayzenberg et al. (2014). Glavan and Lin proposed a prescription leading to an effective four-dimensional Einstein-Gauss-Bonnet (4D-EGB) framework Glavan and Lin (2020), which has been investigated in several directions including consistency analyses Lu and Pang (2020); Hennigar et al. (2020); Gürses et al. (2020), quasinormal modes and shadows Konoplya and Zinhailo (2020); Kumar and Ghosh (2020), and geodesic motion Guo and Li (2020); see Ref. Fernandes et al. (2022) for a review.
@@
Thermodynamic properties of 4D-EGB black holes have been analyzed extensively, including extended thermodynamics and information-geometric diagnostics Wei and Liu (2020); Hosseini Mansoori (2021); Eslam Panah et al. (2020); Hegde et al. (2021, 2024, 2024); Kumar et al. (2022); Kumar and Ghosh (2023); Kumar et al. (2024); Fernandes (2020). Most existing studies of black hole thermodynamics adopt the conventional semiclassical entropy and focus on classical phase structure. For example, Ladghami et al. Ladghami et al. (2023) considered charged black solutions in 4D-EGB gravity and provided a thermodynamic interpretation for GB coupling, in addition to reporting several phase transitions contingent on the interplay between GB coupling and charge of the black holes. Moreover, the possibility of a logarithmic contribution (which is normally reported in quantum gravity paradigm Rovelli (1996); Ashtekar et al. (1998); Kaul and Majumdar (2000); Dabholkar et al. (2011); Mandal and Sen (2010)) beyond the leading Bekenstein-Hawking term to the black hole entropy has been reported by Fernandes Fernandes (2020). Wei and Liu Wei and Liu (2020) applied Ruppeiner information geometry to charged 4D‑EGB black holes, identifying van‑der‑Waals–type small/large black-hole phase behavior, and using the thermodynamic curvature to diagnose regimes suggestive of attractive/repulsive effective interactions. A similar investigation by Mansoori Hosseini Mansoori (2021) utilizes a novel information geometric probe Hosseini Mansoori and Mirza (2019) to quantify the role of GB coupling on black hole microstructures within 4D-EGB gravity. Several other studies that cover various aspects of charged black holes in 4D-EGB gravity can be found in Yang et al. (2020, 2025); Guo et al. (2022); Kumar et al. (2020). Whether the aforementioned thermal behavior in 4D‑EGB gravity changes qualitatively once quantum-scale (short-distance) effects become relevant remains an open question. This motivates us to examine how non-perturbative, short-distance quantum corrections to the entropy reshape stability criteria and information-geometric signatures within the 4D‑EGB landscape.
@@
In this work, we consider a charged 4D-EGB black hole in the GR branch and restrict to the non-extremal regime $M>\sqrt{Q^{2}+\alpha}$ in a fixed-$Q$ (canonical) setup. Quantum effects are incorporated through a non-perturbative exponential deformation of the entropy Chatterjee and Ghosh (2020), while the geometry and the Hawking temperature defined by surface gravity are kept classical, so the results should be interpreted as effective thermodynamic diagnostics Masood A. S. Bukhari et al. (2023). Our main deliverables are: (i) heat capacity and Helmholtz free-energy analyses of stability and phase-transition-like behavior; (ii) an effective quantum-work characterization based on Jarzynski equality; and (iii) the Ruppeiner thermodynamic geometry Ruppeiner (2014) on the $(M,Q)$ state space, where curvature signatures are correlated with heat-capacity structure. The investigations find relevance to extremality and the associated black hole remnants in quantum regime. To our knowledge, this is the first 4D‑EGB information‑geometric study that combines exponential (non‑perturbative) entropy and Jarzynski-based work diagnostics in the near‑extremal regime with relevance to quantum domain.
@@ II Conceptual aspects
where $R$ is the Ricci curvature scalar, and $g$ the determinant of metric tensor $g_{\mu\nu}$. Lovelock theorem Lovelock (1971, 1972); Lanczos (1938) asserts that GR is the unique theory of gravity in four dimensions, provided certain conditions are met. These conditions include diffeomorphism invariance, metricity, and second-order equations of motion. In higher-dimensional spacetime, the action that satisfies these three conditions is the GB action, given by
@@
and $T_{\mu\nu}$ is the stress-energy tensor. Until recently, it was well known that in ordinary 4D spacetime, the fact that $\mathcal{G}$ is a total derivative implies it has no contribution to the dynamics. However, the idea proposed in Ref. Glavan and Lin (2020) suggested a way to circumvent this conclusion through a novel rescaling of the coupling constant by
@@
At the level of the equations of motion, the Glavan–Lin rescaling prescription (6) has been the subject of substantial discussion. In particular, it has been emphasized that taking $D\!\to\!4$ is not a standard continuous limit for tensorial expressions and can lead to ambiguities beyond highly symmetric ansätze Ai (2020), and that the naive construction does not generically define consistent four-dimensional field equations without additional prescriptions Gurses et al. (2020). Related analyses have further argued that, even when linearized fluctuations around maximally symmetric backgrounds are well behaved, the perturbative expansion becomes ill-defined at higher order, indicating that additional structure is required to obtain a well-defined four-dimensional theory Arrechea et al. (2021). Motivated by these issues, several works have proposed explicit $D\!\to\!4$ completions/regularizations—typically involving an additional scalar-type degree of freedom and/or a controlled breaking of diffeomorphism invariance—which aim to provide well-defined four-dimensional field equations Aoki et al. (2020); Lu and Pang (2020); Hennigar et al. (2020); Gürses et al. (2020); Charmousis et al. (2025), and phenomenological/observational implications Charmousis et al. (2022); García-Aspeitia and Hernández-Almada (2021); Wang and Mota (2021); Khodabakhshi et al. (2024). Related discussions and other broader perspectives of the theory can be found in the review work Fernandes et al. (2022). Since there is no unique universally adopted resolution of the dimensional regularization problem, in this work, we employ the spherically symmetric black hole solution arising from the Glavan–Lin approach as an effective phenomenological background and interpret the thermodynamic results within this specific 4D-EGB setting. We adopt this prescription here because it provides a widely used, closed-form static black-hole background that enables direct comparison with earlier thermodynamic studies in the 4D‑EGB literature. In the present work, the rescaling affects our results only through the chosen metric function (and hence the horizon structure/extremality bound), on top of which we implement an exponential entropy deformation while keeping the geometry classical. The conclusions should therefore be interpreted as model-dependent thermodynamic diagnostics within this specific effective 4D‑EGB construction. We thus expect that alternative $D\rightarrow 4$ completions could lead to different quantitative behavior.
@@ II.2 Black holes in 4D-GB theory
The novel theory of 4D-GB gravity thus predicts a line element of the form Glavan and Lin (2020); Fernandes et al. (2022)
@@
which is not a well-defined limit. Note also that the mass term for the GB branch has the wrong sign Glavan and Lin (2020); Fernandes et al. (2022). Hence, this branch is discarded as a viable physical solution. On these grounds, we only consider the GR branch in our analysis.
@@
We graphically illustrate the metric function $f(r)$ and the horizon radius $r_{+}$ against GB parameter $\alpha$ in Fig.1. It can be readily observed from Fig. 1(a) that $f(r)$ for the 4D-GB black hole is finite near the origin $r\rightarrow 0$. However, the metric coefficients being finite at the origin $r\rightarrow 0$ do not eliminate the central singularity, as the Kretschmann curvature scales as $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}\propto r^{-3}$ Fernandes et al. (2022). Interestingly, in GR for a Schwarzschild black hole, the Kretschmann scalar scales as $r^{-6}$ at any radius $r$. This implies that the introduction of the GB coupling parameter $\alpha$ weakens the black hole singularity. One can also see from Fig. 1(b) that for a given black hole mass $M$, the introduction of $\alpha$ reduces the black hole size compared to the GR case. This subtle effect arises purely from higher curvature corrections to Einstein gravity, providing a physical picture of the nontrivial contributions of $\alpha$ to gravitational dynamics. Additionally, we note the presence of another singularity at a radius where $r^{3}=-8\alpha M$, rendering the expression under the square root in Eq. (8) zero. While the parameter $\alpha$ is permitted to take both positive and negative values Guo and Li (2020); Konoplya and Zinhailo (2020), its negative values are subject to stringent constraints Charmousis et al. (2022). In this work, we focus exclusively on positive values of $\alpha$, following the approach in the original study Glavan and Lin (2020).
Figure 2: Black hole temperature $T_{\rm BH}$ vs the mass $M$. The GB coupling $\alpha$ parameter tends to make the black hole colder on smaller scales, mimicking the role of charge in RN geometry.
Charged black hole solutions in 4D-GB gravity have also been found, for which the following relation holds Fernandes (2020)
@@
The quantity $T_{\rm BH}$ is plotted in Fig. 2, illustrating that the GB coupling parameter $\alpha$ reduces the black hole temperature at smaller scales, exhibiting a behavior similar to that of the charge $Q$ in Reissner-Nordström (RN) geometry . Notably, we did not introduce any corrections to $T_{\rm BH}$ due to $\eta$ (see Eq. (18)). The reason for this is that we use a modified entropy formula based on quantum gravity principles Chatterjee and Ghosh (2020) without invoking any quantum geometry. Consequently, the metric function $f(r)$ in Eq. (8) remains unchanged, and the definition of temperature $(T_{\rm BH}\propto\mathrm{d}f(r)/\mathrm{d}r)$ follows from this function. This approach has been discussed in some earlier works Pourhassan et al. (2022a, b); Masood A. S. Bukhari et al. (2023).

### v5

# thanks: Author full legal name (official records): Syed Masood Ahmad Shah Bukhari

@@ Abstract
Glavan and Lin’s proposal of an effective four-dimensional Einstein–Gauss–Bonnet (4D-EGB) gravity framework yields predictions that differ from general relativity in some regimes. A range of black hole studies have offered insights into the dynamical and phenomenological aspects of this effective theory of gravity. In this work, the thermodynamics of a charged 4D-EGB black hole with Gauss–Bonnet (GB) coupling $\alpha$, characterized by mass $M$ and charge $Q$ in the non-extremal regime $M>\sqrt{Q^{2}+\alpha}$ is investigated by combining a non-perturbative, quantum-gravity-inspired exponential correction to the entropy (quantified by $\eta$) with information-geometric diagnostics. Within a canonical ensemble (fixed $Q$) paradigm, thermodynamic stability regions and phase-transition-like features are identified as the black hole size tends toward extremality due to Hawking evaporation. The Ruppeiner metric is then constructed on the $(M,Q)$ state space and the associated thermodynamic curvature is evaluated to characterize the effective interaction signatures and its relation to critical behavior. In addition, an effective quantum-work quantity, defined from the free-energy landscape using Jarzynski equality, is evaluated as an additional probe of short-distance, near-extremal behavior. The results indicate that departures from the general-relativistic behavior are negligible for large black holes but can become relevant at small horizon scales. Specifically, on short-distance scales, the combined influence of $\alpha$ and $\eta$ can modify stability of the extremal black hole geometry and remnants within this thermodynamic model.
@@ I Introduction
Black hole thermodynamics originates from the work of Hawking [49, 50] and Bekenstein [13] in semiclassical gravity, where entropy satisfies the area law,
@@
with $A$ the horizon area and $\ell_{\rm p}$ the Planck length  [^1] In this paper, the natural units $c=G=\hbar=1$ are used throughout. The area scaling is closely related to holography [114, 17]. While this framework provides a consistent macroscopic description in terms of conserved charges, the microscopic origin of black hole entropy and the statistical interpretation of thermodynamic response functions remain open [30, 88].
@@
Hawking evaporation [49, 50] is derived in a semiclassical regime, and motivates corrections to the Bekenstein–Hawking description when the black hole approaches microscopic scales. Many quantum-gravity approaches imply a minimum length and associated departures from classical geometry [60], which in turn suggest modifications of the entropy [83, 115, 94] and connections to microstate counting [113]. Entropy corrections predicted in various frameworks include perturbative (logarithmic or power-law) terms [101, 10, 66, 26, 82] and non-perturbative contributions of exponential type [11, 38, 28, 25]. Examples of the latter arise in AdS/CFT-based analyses [81] and in treatments involving Kloosterman sums [86, 27, 28]. Such corrections are suppressed for macroscopic horizons but can dominate the thermodynamics near microscopic scales [27, 28, 25]; see Refs. [53, 41, 100, 108, 83, 110, 40, 95, 90, 94, 93, 116, 37, 15, 91, 92, 8] for broader discussions.
@@
A complete quantum-gravity treatment would modify both the geometry and thermodynamics of black holes. Nevertheless, it is common in phenomenological studies to implement quantum effects through an effective entropy deformation while keeping the classical background fixed, thereby isolating their impact on response functions and stability criteria [85, 91, 92, 62]. Since thermodynamic behavior depends on the underlying gravitational dynamics, it is also of interest to examine these effects in modified-gravity models.
@@
General relativity (GR) is consistent with a wide range of observations, including gravitational waves [1, 2] and black hole shadow measurements [4], but it also motivates extensions in regimes where curvature becomes large. Higher-curvature theories provide a concrete class of such extensions [18, 14, 111]. Gauss-Bonnet (GB) gravity, originating with Lanczos and Lovelock [77, 78, 79], yields second-order field equations and avoids Ostrogradsky-type instabilities [87], and quadratic curvature terms appear naturally in string-inspired effective actions [125, 43, 42]. In four dimensions the GB invariant does not contribute to the classical field equations unless coupled to additional fields, such as scalars [16, 69, 84, 12]. Glavan and Lin proposed a prescription leading to an effective four-dimensional Einstein-Gauss-Bonnet (4D-EGB) framework [39], which has been investigated in several directions including consistency analyses [80, 56, 47], quasinormal modes and shadows [68, 74], and geodesic motion [44]; see Ref. [34] for a review.
@@
Thermodynamic properties of 4D-EGB black holes have been analyzed extensively, including extended thermodynamics and information-geometric diagnostics [120, 59, 31, 52, 51, 51, 71, 73, 72, 35]. Most existing studies of black hole thermodynamics adopt the conventional semiclassical entropy and focus on classical phase structure. For example, Ladghami et al. [76] considered charged black solutions in 4D-EGB gravity and provided a thermodynamic interpretation for GB coupling, in addition to reporting several phase transitions contingent on the interplay between GB coupling and charge of the black holes. Moreover, the possibility of a logarithmic contribution (which is normally reported in quantum gravity paradigm [101, 10, 66, 26, 82]) beyond the leading Bekenstein-Hawking term to the black hole entropy has been reported by Fernandes [35]. Wei and Liu [120] applied Ruppeiner information geometry to charged 4D‑EGB black holes, identifying van‑der‑Waals–type small/large black hole phase behavior, and using the thermodynamic curvature to diagnose regimes suggestive of attractive/repulsive effective interactions. A similar investigation by Mansoori [59] utilizes a novel information geometric probe [58] to quantify the role of GB coupling on black hole microstructures within 4D-EGB gravity. Several other studies that cover various aspects of charged black holes in 4D-EGB gravity can be found in refs.[123, 124, 45, 75]. Whether the aforementioned thermal behavior in 4D‑EGB gravity changes qualitatively once quantum-scale (short-distance) effects become relevant remains an open question. This motivates us to examine how non-perturbative, short-distance quantum corrections to the entropy reshape stability criteria and information-geometric signatures within the 4D‑EGB landscape.
@@
In this work, we consider a charged 4D-EGB black hole in the GR branch and restrict to the non-extremal regime $M>\sqrt{Q^{2}+\alpha}$ in a fixed-$Q$ (canonical) setup. Quantum effects are incorporated through a non-perturbative exponential deformation of the entropy [25], while the geometry and the Hawking temperature defined by surface gravity are kept classical, so the results should be interpreted as effective thermodynamic diagnostics [85]. Our main deliverables are: (i) heat capacity and Helmholtz free-energy analyses of stability and phase-transition-like behavior; (ii) an effective quantum-work characterization based on Jarzynski equality; and (iii) the Ruppeiner thermodynamic geometry [106] on the $(M,Q)$ state space, where curvature signatures are correlated with heat-capacity structure. The investigations find relevance to extremality and the associated black hole remnants in quantum regime. To our knowledge, this is the first 4D‑EGB information‑geometric study that combines exponential (non‑perturbative) entropy and Jarzynski-based work diagnostics in the near‑extremal regime with relevance to quantum domain.
@@ II Conceptual aspects
where $R$ is the Ricci curvature scalar, and $g$ the determinant of metric tensor $g_{\mu\nu}$. Lovelock theorem [78, 79, 77] asserts that GR is the unique theory of gravity in four dimensions, provided certain conditions are met. These conditions include diffeomorphism invariance, metricity, and second-order equations of motion. In higher-dimensional spacetime, the action that satisfies these three conditions is the GB action, given by
@@
and $T_{\mu\nu}$ is the stress-energy tensor. Until recently, it was well known that in ordinary 4D spacetime, the fact that $\mathcal{G}$ is a total derivative implies it has no contribution to the dynamics. However, the idea proposed in Ref. [39] suggested a way to circumvent this conclusion through a novel rescaling of the coupling constant by
@@
At the level of the equations of motion, the Glavan–Lin rescaling prescription (6) has been the subject of substantial discussion. In particular, it has been emphasized that taking $D\!\to\!4$ is not a standard continuous limit for tensorial expressions and can lead to ambiguities beyond highly symmetric ansätze [3], and that the naive construction does not generically define consistent four-dimensional field equations without additional prescriptions [46]. Related analyses have further argued that, even when linearized fluctuations around maximally symmetric backgrounds are well behaved, the perturbative expansion becomes ill-defined at higher order, indicating that additional structure is required to obtain a well-defined four-dimensional theory [9]. Motivated by these issues, several works have proposed explicit $D\!\to\!4$ completions/regularizations—typically involving an additional scalar-type degree of freedom and/or a controlled breaking of diffeomorphism invariance—which aim to provide well-defined four-dimensional field equations [7, 80, 56, 47, 23], and phenomenological/observational implications [24, 36, 117, 67]. Related discussions and other broader perspectives of the theory can be found in the review work [34]. Since there is no unique universally adopted resolution of the dimensional regularization problem, in this work, we employ the spherically symmetric black hole solution arising from the Glavan–Lin approach as an effective phenomenological background and interpret the thermodynamic results within this specific 4D-EGB setting. We adopt this prescription here because it provides a widely used, closed-form static black hole background that enables direct comparison with earlier thermodynamic studies in the 4D‑EGB literature. In the present work, the rescaling affects our results only through the chosen metric function (and hence the horizon structure/extremality bound), on top of which we implement an exponential entropy deformation while keeping the geometry classical. The conclusions should therefore be interpreted as model-dependent thermodynamic diagnostics within this specific effective 4D‑EGB construction. We thus expect that alternative $D\rightarrow 4$ completions could lead to different quantitative behavior.
@@ II.2 Black holes in 4D-GB theory
The novel theory of 4D-GB gravity thus predicts a line element of the form [39, 34]
@@
which is not a well-defined limit. Note also that the mass term for the GB branch has the wrong sign [39, 34]. Hence, this branch is discarded as a viable physical solution. On these grounds, we only consider the GR branch in our analysis.
@@
We graphically illustrate the metric function $f(r)$ and the horizon radius $r_{+}$ against GB parameter $\alpha$ in Fig.1. It can be readily observed from Fig. 1(a) that $f(r)$ for the 4D-GB black hole is finite near the origin $r\rightarrow 0$. However, the metric coefficients being finite at the origin $r\rightarrow 0$ do not eliminate the central singularity, as the Kretschmann curvature scales as $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}\propto r^{-3}$ [34]. Interestingly, in GR for a Schwarzschild black hole, the Kretschmann scalar scales as $r^{-6}$ at any radius $r$. This implies that the introduction of the GB coupling parameter $\alpha$ weakens the black hole singularity. One can also see from Fig. 1(b) that for a given black hole mass $M$, the introduction of $\alpha$ reduces the black hole size compared to the GR case. This subtle effect arises purely from higher curvature corrections to Einstein gravity, providing a physical picture of the nontrivial contributions of $\alpha$ to gravitational dynamics. Additionally, we note the presence of another singularity at a radius where $r^{3}=-8\alpha M$, rendering the expression under the square root in Eq. (8) zero. While the parameter $\alpha$ is permitted to take both positive and negative values [44, 68], its negative values are subject to stringent constraints [24]. In this work, we focus exclusively on positive values of $\alpha$, following the approach in the original study [39].
Figure 2: Black hole temperature $T_{\rm BH}$ versus the mass $M$. The GB coupling $\alpha$ parameter tends to make the black hole colder on smaller scales, mimicking the role of charge in RN geometry.
Charged black hole solutions in 4D-GB gravity have also been found, for which the following relation holds [35]
@@
The quantity $T_{\rm BH}$ is plotted in Fig. 2, illustrating that the GB coupling parameter $\alpha$ reduces the black hole temperature at smaller scales, exhibiting a behavior similar to that of the charge $Q$ in Reissner-Nordström (RN) geometry . Notably, we did not introduce any corrections to $T_{\rm BH}$ due to $\eta$ (see Eq. (18)). The reason for this is that we use a modified entropy formula based on quantum gravity principles [25] without invoking any quantum geometry. Consequently, the metric function $f(r)$ in Eq. (8) remains unchanged, and the definition of temperature $(T_{\rm BH}\propto\mathrm{d}f(r)/\mathrm{d}r)$ follows from this function. This approach has been discussed in some earlier works [91, 92, 85].